MILLIMETER-WAVE SPECTROSCOPY OF CoNO IN THE GROUND (X^{1}Σ) STATE

AI SAKAMOTO, MASATO HAYASHI, \underline{KENSUKE HARADA,} KEIICHI TANAKA and TAKE-HIKO TANAKA, Department of Chemistry, Faculty of Sciences, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN.

Infrared spectrum of CoNO has been measured in low temperature Ar matrix and the ν_{1} and ν_{3} band origins were reported to be 1761.0 and 620.1 cm^{−1} respectively. ^{a} Recent DFT calculations suggested that the linear form (X^{1}Σ) and bent form (X^{3}A′) of CoNO have almost the same energies.^{a} In the present study, the pure rotational spectrum of CoNO generated in a supersonic jet expansion by ultraviolet photolysis of Co(CO)_{3}NO was observed in the millimeter-wave region. This is the first observation of the rotationally resolved spectrum of the transition metal nitrosyl in the gas phase. Seven rotational transitions (from J = 6 − 5 to 12 − 11) were measured in the frequency region of 56 − 112 GHz. Each rotational transition was split into 8−11 components due to hyperfine interaction of the Co (I = 7/2) nucleus. The spectrum was analyzed to determine molecular constants, including the rotational constant B, centrifugal distortion constant D, nuclear quadrupole interaction constant eQq, and nuclear spin-rotation interaction constant C_{I}. From the observed spectral pattern, it is confirmed that CoNO has a linear structure and the electronic ground state is ^{1}Σ. The Co−N bond length was calculated to be 1.588 Å from the rotational constant, which is by 0.1 Å shorter than the Co−C bond length of CoCO. Measurement of rotational transitions in the ν_{2} vibrationally excited state is in progress.

aMingfei Zhou and Lester Andrews, J. Phys. Chem., A104, 3915 (2000)